# PoLi-RL: A Point-to-List Reinforcement Learning Framework for Conditional Semantic Textual Similarity

### v1

@@ 2.3 PoLi-RL: A Two-Stage Reinforcement Learning Framework
As previously established, our optimization goal is to maximize the expected reward, making the design of the reward function $\mathcal{R}$ central to our method. To ensure stable and effective optimization, we propose PoLi-RL, a framework built upon a two-stage progressive reward curriculum. This subsection details our pipeline and its reward mechanisms, as illustrated in Figure 2.
Figure 2: An overview of the PoLi-RL framework. The framework employs a two-stage curriculum, progressing from Stage I, where the model learns foundational scoring rules, to Stage II, which refines the model’s ability to discern fine-grained semantic differences. The core of our method is the PSRR mechanism in Stage II, where pairwise and listwise ranking rewards are computed vertically within slices of same-indexed completions to provide precise, differentiated learning signals.
@@
The goal of Stage I is to ground the model in the fundamental scoring rules of the C-STS task. For each input sample, the policy generates $G$ completions, from which we parse a set of predicted scores $\{\tilde{y}_{j}\}_{j=1}^{G}$. The total reward for Stage I, $R_{S1}$, is a weighted sum of three components:
@@
The Pointwise Accuracy Reward ($R_{\text{pointwise}}$) is the primary component in Stage I. It measures the normalized distance between the predicted score $\tilde{y}_{j}$ and the ground-truth score $y_{j}$.
@@
To counter reward hacking, where the model tends to output safe intermediate scores, we introduce a Binary Judgement Reward ($R_{\text{binary}}$). According to the C-STS guideline that scores $\geq 3$ indicate similarity while scores $\leq 2$ indicate dissimilarity, this reward encourages the model to first master this basic binary classification:
@@
Finally, a simple Format Reward ($R_{\text{format}}$) ensures the output adheres to the required structure, which consists of a binary judgment (‘yes’ or ‘no’) followed by the numerical score in parentheses.
Stage II: Fine-Grained Semantic Refinement.
After the model has acquired basic scoring abilities in Stage I, Stage II refines its capacity to discern subtle semantic differences by introducing a richer, hybrid reward signal that incorporates both pairwise and listwise ranking objectives.
PSRR: A Fine-Grained Reward Mechanism.
A primary challenge in directly optimizing ranking metrics is that a single, batch-wide reward signal is too coarse to assign credit precisely. To address this, we propose the Parallel Slice Ranking Reward (PSRR) mechanism. The core idea of PSRR is to restructure the generated outputs to enable more granular reward computation. For a batch of $N$ samples, we begin by having the policy generate $G$ completions $\{o_{i,1},\dots,o_{i,G}\}$ for each sample $x_{i}$. From each completion $o_{i,j}$, we then parse the predicted score $\tilde{y}_{i,j}$. Instead of treating the completions as a single flat list, we organize these $N\times G$ predicted scores into $G$ “parallel slices”. Each slice, denoted as $Y^{j}_{pred}$, is defined as the collection of the $j$-th predicted score from all $N$ samples in the batch: $Y^{j}_{pred}=\{\tilde{y}_{1,j},\tilde{y}_{2,j},\ldots,\tilde{y}_{N,j}\}$, where $j\in\{1,\ldots,G\}$. This slicing architecture is the foundation upon which our advanced ranking rewards are built, ensuring that each completion receives a specific learning signal based on its relative performance within its slice.
A sufficiently large slice size $N$ is crucial for a stable and meaningful ranking signal. To make this computationally feasible with limited GPU memory, we leverage gradient accumulation. Specifically, we first generate the full set of $N\times G$ completions across multiple devices and accumulation steps. These completions are then organized into parallel slices to compute rewards and advantages globally. Subsequently, for the forward and backward passes that compute loss and gradients, we process smaller sub-batches sequentially, accumulating gradients over multiple passes before executing a single optimizer step. This strategy makes our reward design practically feasible, enabling the model to learn from a rich, large-scale ranking signal without requiring prohibitive memory.
@@
Computed within each parallel slice $P_{j}$, this reward leverages the paired structure of the C-STS dataset to provide a local ranking signal. It is applied only to adjacent input samples ($x_{i},x_{i+1}$) that form a valid pair. For such a pair, we define the predicted difference as $\Delta_{\text{pred}}=\tilde{y}_{i,j}-\tilde{y}_{i+1,j}$ and the true difference as $\Delta_{\text{true}}=y_{i}-y_{i+1}$. The reward $R^{\text{pairwise}}_{i,j}$ is then a piecewise function:
$R^{\text{pairwise}}_{i,j}=\begin{cases}0&\text{if }\text{sgn}(\Delta_{\text{pred}})\neq\text{sgn}(\Delta_{\text{true}})\\ R_{\text{base}}+(1-R_{\text{base}})\cdot\left(1-\frac{\left|\Delta_{\text{pred}}-\Delta_{\text{true}}\right|}{\text{max\_error}}\right)&\text{if }\text{sgn}(\Delta_{\text{pred}})=\text{sgn}(\Delta_{\text{true}})\end{cases}$ (7)
This function first checks if the basic ranking preference is correct using the sign function $\text{sgn}(\cdot)$. If the order is wrong, the reward is zero. If correct, a base reward $R_{\text{base}}$ is given, plus an additional reward that measures the normalized distance between the predicted and true score differences. Here, max_error is the maximum possible score difference, which is 3 for paired samples.
@@
While the pairwise reward focuses on local comparisons, the listwise reward provides a more global ranking perspective within each slice. It is calculated as the normalized difference between a completion’s predicted rank within its slice and the ideal rank of its ground-truth label, formulated as:
$R^{\text{listwise}}_{i,j}=1-\frac{|\text{Rank}(\tilde{y}_{i,j},Y^{j}_{pred})-\text{Rank}(y_{i},Y_{\text{true}})|}{N-1}$ (8)
where $Y_{\text{true}}=\{y_{1},\ldots,y_{N}\}$ is the set of true labels for the current batch, the function $\text{Rank}(v,S)$ returns the rank of element $v$ within the set $S$ (from 1 to $N$), and the division by N1N-1 normalizes the rank error to the range $[0,1]$.
The final reward for Stage II, $R_{S2}$, combines the robust Pointwise Reward from Stage I as a stabilizing anchor with the new ranking-based rewards enabled by PSRR. The total reward is a weighted combination of these three components:
@@ 3 Experiments
To empirically validate the effectiveness of our proposed PoLi-RL framework, we conduct a comprehensive set of experiments. We begin by detailing our experimental setup in subsection 3.1, including the dataset, evaluation metrics, baselines, and implementation details. Following this, in subsection 3.2, we present the main results, comparing PoLi-RL against a suite of strong baselines. Finally, in subsection 3.3, we conduct a series of ablation studies to analyze the contributions of our framework’s key components.
@@ 3.1 Experimental Setup
We build PoLi-RL upon the Qwen3-8B (Yang et al., 2025) model using the ms-swift framework (Zhao et al., 2024) for RL training. All experiments are conducted on the official C-STS dataset (Deshpande et al., 2023). Following prior work, we use Spearman correlation as the primary metric and Pearson as the secondary. We compare our method against three baseline categories: First, the discriminative models in the cross-encoder setting. Second, powerful generative LLMs, such as Flan-UL (Tay et al., 2022), Flan-T5 (Chung et al., 2024) and Tk-Instruct (Wang et al., 2022). Finally, our own SFT and few-shot implementations on Qwen3-8B for direct comparison.
@@
PoLi-RL (Ours) RL 8B 48.18 48.27
3.2 Main Results
Table 1 summarizes the performance of our framework. PoLi-RL achieves Spearman and Pearson correlation coefficients of 48.18 and 48.27, respectively, establishing a new state-of-the-art (SOTA) for the cross-encoder architecture.
Comparing against prior work and different training paradigms reveals several key insights. First, PoLi-RL surpasses the previous cross-encoder SOTA, SEAVER (Li et al., 2024), by a significant margin of 4.35 points. Second, when compared to our own implementations on the same Qwen3-8B model, PoLi-RL achieves substantial improvements of 7.76 points over SFT and 10.28 points over few-shot prompting. The advantage over SFT is particularly noteworthy: while supervised fine-tuning relies on token-level cross-entropy loss that only implicitly encourages correct outputs, reinforcement learning enables explicit optimization through carefully designed reward signals that can directly guide the model’s reasoning process and better capture the nuanced requirements of conditional similarity judgment.
[…]
@@ 3.3 Ablation Studies
To understand the source of PoLi-RL’s performance gains and validate our design choices, we conduct a series of ablation studies.
@@
Table 2 dissects the effectiveness of our progressive training schedule. We first observe that a Naive RL approach (Row 2), which uses only a single, batch-wise listwise reward from scratch, yields negligible improvement over the few-shot baseline (Row 1), demonstrating the need for a more structured curriculum. Our PoLi-RL Stage I (Row 3) addresses this by building a robust foundation, substantially outperforming the few-shot baseline by 6.87 points. Within this stage, ablating the binary reward (Row 4) leads to a discernible dip in performance, validating its role in grounding the model in the task’s basic binary judgment.
@@
Sensitivity to Reward Weights in PoLi-RL Stage II.
Table 3 details our analysis of the framework’s sensitivity to reward weights in Stage II. The results reveal a delicate balance, with peak performance achieved by moderately increasing the pairwise weight to 1.5. The model also shows considerable tolerance to pairwise signal, as halving its weight to 0.5 results in only a marginal performance drop to 47.77. This suggests that the stable, local ranking signal from the pairwise reward is instrumental in refining the model’s performance to its optimal state. In contrast, the framework is more sensitive to the pointwise and listwise weights, where deviations from their baseline of 1.0 tend to hinder performance. Crucially, the framework’s performance remains robust across all reward weight configurations. This demonstrates that our hybrid reward, anchored by the pointwise signal and refined by the PSRR mechanism, provides a consistently effective learning signal.
@@
Sensitivity to Parallel-Slice Size.
To determine the optimal configuration for our PSRR mechanism, we study the impact of the slice size $N$, with results presented in Table 4. The empirical results reveal a clear trend: performance peaks at an intermediate slice size of $N=24$ and degrades as the size deviates in either direction. This suggests that an optimal balance is required for the ranking signal. A slice that is too small may provide a less stable ranking signal, while one that is too large makes the ranking task overly complex for the model to learn effectively. This finding validates the design principle behind our PSRR mechanism: a carefully-sized, localized ranking signal is more effective than a purely global or an overly-restricted one.
@@ 4 Analysis
In this section, we provide a quantitative and qualitative analysis of PoLi-RL’s performance, focusing on its error distribution and nuanced reasoning capabilities.
@@ 4.1 Analysis of Prediction Error Distribution
Figure 3 visualizes the distribution of absolute prediction errors on the C-STS validation set. The plot reveals several key insights into the models’ behaviors. While an error of 1 is the most frequent outcome for all models, likely reflecting the inherent nuances of the C-STS scale, a clear progression of improvement is evident. Compared to the raw and SFT model, PoLi-RL demonstrates a more favorable error distribution. Firstly, it achieves the highest density of perfect predictions (error=0). More importantly, the density of the PoLi-RL curve in the high-error regions is consistently the lowest. This shows that our method significantly reduces the frequency of large, unreliable errors, yielding a more stable and reliable model.

### v2

@@ 2.3 PoLi-RL: A Two-Stage Reinforcement Learning Framework
To optimize effectively, we propose PoLi-RL, a framework built on a two-stage progressive reward curriculum. This subsection elaborates our pipeline and its reward mechanisms (Figure 2).
Figure 2: An overview of the PoLi-RL framework. It employs a two-stage curriculum, progressing from Stage I, where the model learns foundational scoring rules, to Stage II, which refines the model’s ability to discern fine-grained semantic differences. The core of our method is the PSRR mechanism in Stage II, where pairwise and listwise ranking rewards are computed vertically within slices of the same-indexed completions to provide precise, differentiated learning signals.
@@
The goal of Stage I is to ground the model in the fundamental scoring rules of the C-STS task. For each input sample, the LLM policy generates $G$ completions, from which we parse a set of predicted scores $\{\tilde{y}_{j}\}$. The total reward for Stage I, $R_{S1}$, is a weighted sum of three components:
@@
The Pointwise Accuracy Reward ($R_{\text{pointwise}}$) is the primary component in Stage I. It measures the normalized distance between the predicted score $\tilde{y}_{j}$ and the ground-truth score $y_{j}$:
@@
To mitigate the model’s tendency to converge to safe, intermediate scores, we introduce a Binary Judgment Reward ($R_{\text{binary}}$). According to the C-STS guideline that scores $\geq 3$ indicate similarity, while scores $\leq 2$ indicate dissimilarity, this reward encourages the model to first master this basic binary classification:
@@
Finally, a simple Format Reward ($R_{\text{format}}$) ensures the output adheres to the required structure, which consists of a binary judgment (“yes” or “no”) followed by the numerical score in parentheses.
Stage II: Fine-Grained Semantic Distinction.
Building on the foundational scoring abilities from Stage I, Stage II refines the model’s capacity for nuanced semantic distinctions. This is achieved through a hybrid reward signal that incorporates both pairwise and listwise ranking objectives, both of which are enabled by our proposed Parallel Slice Ranking Reward (PSRR) mechanism.
Parallel Slice Ranking Reward Mechanism.
A primary challenge in directly optimizing ranking metrics is that a single, batch-level reward signal is too coarse to assign credit precisely. To address this, we propose the Parallel Slice Ranking Reward (PSRR) mechanism. The core idea of PSRR is to restructure the generated outputs to enable more granular reward computation. For a batch of $N$ samples, we begin by having the policy generate $G$ completions $\{o_{i,1},\dots,o_{i,G}\}$ for each sample. From each completion $o_{i,j}$, we then parse the predicted score $\tilde{y}_{i,j}$. Instead of treating the completions as a single flat list, we organize these $N\times G$ predicted scores into $G$ “parallel slices”. Each slice, denoted as $Y^{j}_{\text{pred}}$, is defined as the collection of the $j$-th predicted score from all $N$ samples in the batch: $Y^{j}_{\text{pred}}=\{\tilde{y}_{1,j},\tilde{y}_{2,j},\ldots,\tilde{y}_{N,j}\}$, where $j\in\{1,\ldots,G\}$. This slicing architecture forms the foundation of our ranking rewards, ensuring each completion receives a specific learning signal based on its relative performance within its independent slice.
A sufficiently large batch size $N$ is crucial for a stable and meaningful ranking signal. To make this computationally feasible with limited GPU memory, we employ gradient accumulation strategy. Specifically, we first generate the full set of $N\times G$ completions and organize them into parallel slices to compute rewards and advantages globally. Subsequently, we process smaller sub-batches sequentially to compute losses and accumulate gradients over multiple backward passes before executing a single optimizer step. This strategy enables the model to learn from a rich ranking signal without incurring prohibitive memory costs.
@@
Computed within each parallel slice $Y^{j}_{\text{pred}}$, this reward leverages the paired structure of the C-STS dataset to provide a local ranking signal. It is applied only to adjacent input samples that form a valid pair. We define the predicted difference as $\Delta_{\text{pred}}=\tilde{y}_{i,j}-\tilde{y}_{i+1,j}$ and the true difference as $\Delta_{\text{true}}=y_{i}-y_{i+1}$. To accommodate cases where paired samples share identical ground-truth scores ($\Delta_{\text{true}}=0$), we distinguish them from strictly ordered pairs. The reward $R^{\text{pairwise}}_{i,j}$ is formulated as:
$R^{\text{pairwise}}_{i,j}=\begin{cases}1-\frac{|\Delta_{\text{pred}}|}{D_{\max}}&\text{if }\Delta_{\text{true}}=0\\ R_{\text{base}}+(1-R_{\text{base}})\cdot\left(1-\frac{\left|\Delta_{\text{pred}}-\Delta_{\text{true}}\right|}{D_{\max}}\right)&\text{if }\text{sgn}(\Delta_{\text{pred}})=\text{sgn}(\Delta_{\text{true}})\\ 0&\text{otherwise}\end{cases}$ (7)
where $D_{\max}$ represents the maximum possible score difference (i.e., 4 for the 1-5 scale). For strictly ordered pairs ($\Delta_{\text{true}}\neq 0$), the model receives a base reward $R_{\text{base}}$ for correct ordinality, with an additional penalty for prediction error magnitude. For tied pairs ($\Delta_{\text{true}}=0$), the reward simply minimizes the absolute prediction difference.
@@
While the pairwise reward focuses on local comparisons, the listwise reward provides a more global ranking perspective within each slice. It is calculated as the normalized difference between a completion’s predicted rank and its ground-truth rank:
$R^{\text{listwise}}_{i,j}=1-\frac{|\text{Rank}(\tilde{y}_{i,j},Y^{j}_{\text{pred}})-\text{Rank}(y_{i},Y_{\text{true}})|}{N-1}$ (8)
where $Y_{\text{true}}=\{y_{1},\ldots,y_{N}\}$ is the set of ground-truth labels for the current batch, the function $\text{Rank}(v,S)$ returns the rank of element $v$ within the set $S$ (from 1 to $N$), and the division by N1N-1 normalizes the rank error to the range $[0,1]$.
The final reward for Stage II, $R_{S2}$, combines the robust Pointwise Reward from Stage I as an anchor with the ranking-based rewards enabled by PSRR. The total reward is a weighted combination of these three components:
@@ 3.1 Experimental Setup
We build PoLi-RL upon the Qwen3 (Yang et al., 2025) model family (0.6B, 4B, and 8B) using the ms-swift framework (Zhao et al., 2024) for RL training. Detailed hyperparameter configurations are provided in Appendix A.3. All experiments are conducted on the official C-STS dataset (Deshpande et al., 2023). Following prior work, we use Spearman correlation as the primary metric and Pearson as the secondary. We compare our method against three baseline categories: First, discriminative models in the cross-encoder setting, such as SimCSE (Gao et al., 2022). Second, state-of-the-art proprietary and open-source reasoning models, including Deepseek-R1 (Guo et al., 2025), GPT-4 (Achiam et al., 2023), GPT-4o, Flan-T5 (Chung et al., 2024) , Tk-Instruct (Wang et al., 2022), etc. Finally, our own SFT and few-shot implementations on Qwen3 backbones for direct comparison.
3.2 Main Results
Table 1 summarizes the performance of PoLi-RL. Our framework establishes a new state-of-the-art (SOTA) for the cross-encoder architecture, with the 8B model achieving a Spearman correlation of 48.18. The significance of this achievement is underscored by several key observations:
Superiority over Strong Baselines. PoLi-RL significantly outperforms the previous cross-encoder SOTA, SEAVER (Li et al., 2024), by a margin of 4.35 points. Compared to standard SFT, our method yields consistent gains across all model sizes. Notably, on Qwen3-0.6B, it yields substantial absolute improvements of 19.09 points over few-shot inference and 8.75 points over standard SFT, showcasing the substantial benefits of our progressive, multi-component reward optimization.
Efficacy Across Model Scales. A striking finding is the efficacy of PoLi-RL on smaller architectures. As shown in the results for Qwen3-0.6B and 4B, our framework enables compact models to perform significantly beyond their scale. Remarkably, our 0.6B model (44.34) not only outperforms the proprietary giant GPT-4 (43.6) but also surpasses the previous SOTA SEAVER (43.83). This demonstrates that the performance gains stem from aligning the reasoning process with ranking objectives, rather than relying solely on parameter scaling.
Advantage over Proprietary Models. Benchmarking against state-of-the-art reasoning models, our 8B model outperforms GPT-4o by 3.95 points and DeepSeek-R1 by 5.33 points. This suggests that while proprietary models possess strong general reasoning capabilities, they struggle to strictly align with the fine-grained quantization standards of C-STS in a few-shot setting. PoLi-RL bridges this gap by explicitly optimizing this alignment via RL, proving that a specialized, smaller model can surpass general-purpose giants on complex conditional ranking tasks.
Beyond standard benchmarks, we conduct rigorous validation to ensure the robustness of our claims. Comparisons with differentiable ranking objectives (Appendix A.4) confirm the superiority of RL over regression-based approaches. Furthermore, evaluation on the re-annotated C-STS dataset (Appendix A.5) validates PoLi-RL’s SOTA performance, demonstrating that its gains stem from genuine reasoning improvements rather than overfitting to label noise.
@@ 3.2 Main Results
GPT-4o Few-shot - 44.23 44.07
Our Implementation on Qwen3-0.6B
Qwen3-0.6B Few-shot 0.6B 25.25 25.19
Qwen3-0.6B SFT 0.6B 35.59 36.83
PoLi-RL (Qwen3-0.6B) RL 0.6B 44.34 44.36
Our Implementation on Qwen3-4B
Qwen3-4B Few-shot 4B 37.97 38.48
Qwen3-4B SFT 4B 38.41 39.45
PoLi-RL (Qwen3-4B) RL 4B 46.23 46.19
@@
PoLi-RL (Qwen3-8B) RL 8B 48.18 48.27
@@ 3.3 Ablation Studies
Table 2 dissects the effectiveness of our progressive training schedule. We observe that a Naive RL approach (Row 2), which relies solely on a batch-level listwise objective, yields negligible improvement over the few-shot baseline (Row 1), highlighting the necessity of the progressive curriculum. Our PoLi-RL Stage I (Row 3) addresses this by building a robust foundation, substantially outperforming the few-shot baseline by 6.87 points. Within this stage, ablating the binary reward (Row 4) leads to a discernible dip in performance, validating its role in grounding the model in the task’s basic binary judgment.
@@
Sensitivity to Hyperparameters.
We analyze the framework’s robustness to key hyperparameters across both training stages. Regarding reward weights, Table 3 shows that Stage II is robust to weight variations, with peak performance achieved by moderately increasing the pairwise weight to 1.5. The model performs strongly even with the default equidistant weights (1:1:1), and shows considerable tolerance to pairwise signal, as halving its weight to 0.5 results in only a marginal performance drop to 47.77. Similarly, for the pointwise and listwise weights, deviations from their baseline of 1.0 result in only minor fluctuations. Crucially, the framework exhibits stable convergence across all configurations without training collapse. This confirms that our hybrid reward design effectively resolves the optimization difficulties encountered by naive listwise approaches.
@@
To determine the optimal configuration for our PSRR mechanism, we study the impact of the slice size $N$, with results presented in Table 4. The empirical results reveal a clear trend: performance peaks at an intermediate slice size of $N=24$ and degrades as the size deviates in either direction. This suggests that an optimal balance is required for the ranking signal. A slice that is too small may provide a less stable ranking signal, while one that is too large makes the ranking task overly complex for the model to learn effectively. This finding validates the design principle behind our PSRR mechanism: a carefully-sized, localized ranking signal is more effective than a purely global or an overly-restricted one. Further sensitivity analyses on the binary reward weights, pairwise base reward ($R_{\text{base}}$), and generation multiplicity ($G$) are provided in Appendix A.3, consistently validating the robustness of our default configurations.
@@
3.4 Generalizability of PSRR Mechanism
To empirically validate the generalizability of our framework beyond the C-STS domain, we extend our evaluation of PoLi-RL to the WMT-QE 2020 task (Fomicheva et al., 2020). While this task shares the overarching goal of optimizing global Spearman correlation, it diverges fundamentally from C-STS across three dimensions: (1) Domain: multilingual translation quality estimation, (2) Label Scale: continuous scores from 0 to 100, and (3) Data Structure: independent samples lacking the deterministic paired ordinality of C-STS.
For this experiment, we employ a minimalist configuration of PoLi-RL by disabling the C-STS-specific pairwise and binary rewards. We relied solely on the pointwise reward (Stage I) and the PSRR-based listwise ranking reward (Stage II), with prompt settings adapted from Sato et al. (2024).
As shown in Table 5, even with this simplified setup, PoLi-RL achieves a Spearman correlation of 54.33, yielding a substantial 3.43 point gain over the strong SFT baseline (50.90). Furthermore, the progression from Stage I (51.72) to Stage II explicitly highlights the added value of the listwise ranking objective. These results confirm that the PSRR mechanism is not merely a task-specific optimization for C-STS, but a transferable solution for aligning LLMs with listwise ranking objectives across domains and label scales.
Table 5: Generalization performance on WMT-QE 2020 task (en-zh subset) using only the core PSRR mechanism, excluding C-STS-specific reward components.
Methods Training Paradigm Spearman $\uparrow$ Pearson $\uparrow$
Qwen3-8B Few-shot 45.03 44.18
Qwen3-8B SFT 50.90 51.09
PoLi-RL (Stage I) RL 51.72 50.58
PoLi-RL (Stage II) RL 54.33 51.09
@@ 4.1 Analysis of Prediction Error Distribution
Figure 3 visualizes the distribution of absolute prediction errors on the C-STS validation set. The plot reveals several key insights into the models’ behaviors. While an error of 1 is the most frequent outcome for all models, likely reflecting the inherent nuances of the C-STS scale, a clear progression of improvement is evident. Compared to the few-shot and SFT model, PoLi-RL demonstrates a more favorable error distribution. It not only achieves the highest density of perfect predictions (error=0) but, more importantly, effectively suppresses the probability of high-error predictions (errors $\geq$ 2). This indicates that our method significantly reduces the frequency of large, unreliable deviations, yielding a model that is more precise against severe misjudgments.